OpenStudy (anonymous):

Three students are running for class president at Madison High School. Helen has h votes, Jami has h - 34 votes, and Rick has h + 14 votes. Which of the following statements is true regarding Rick's votes? Rick has 14 more votes than Jami. Rick has 48 more votes than Jami. Rick has 14 less votes than Helen. Rick has 48 more votes than Helen.

OpenStudy (anonymous):

@uri

OpenStudy (uri):

brb helping my little bro with his art work.

OpenStudy (uri):

Helen =h Jammie =h-34 Rick =h+14 Make an equation and solve.

OpenStudy (anonymous):

Which of the following interpretations for the given expression is correct? x+y/2 The quotient of x + y and 2. The sum of and y. The sum of x and . The product of the quotient of x and 2 and .